OpenStudy (anonymous):

This year the boys' baseball team won 4 times as many games as it lost. The team played a total of 40 games. How many games did the team win?

OpenStudy (btaylor):

for a= games lost and w = games won, 4a = w --> a = 0.25w a + w = 40 0.25w + w = 40 1.25w = 40 w = 40/1.25 = 32

OpenStudy (anonymous):

Thank you!(: